A student attempted to generate an equivalent expression using the distributive property, as shown below:
2(x + 5) = 2x + 5
What was the mistake made?
A. 2 was not multiplied by x
B. 2 was not multiplied by 5
C. Incorrect sign was used for the first term
D. Incorrect sign was used for the second term

Answer by Theo(13342)

with the distributive law, every term in the first multiplier needs to multiply every term in the second multiplier.

a * (b + c) = (a * b) + (a * c)

(a + b) * (c + d) = (a * c) + (a * d) + (b * c) + (b * d).

in this problem:

2 * (x + 5) should be (2 * x) + (2 * 5).

selection B is your solution.